xpq+yq^2=1 using charpits method


1
Expert's answer
2021-01-19T14:36:23-0500

"xpq+yq^2=1\\
xpq+yq^2-1=0.......(1)\\"

Compare (1) to "f(x,y,z,p,q)=0"

"f(x,y,z,p,q)=xpq+yq^2-1=0.......(2)"

Now, the Charpit's auxiliary equation is:

"\frac{dx}{-\frac{\partial f}{\partial p}}=\frac{dy}{-\frac{\partial f}{\partial q}}=\frac{dz}{-p\frac{\partial f}{\partial p}-q\frac{\partial f}{\partial p}}=\frac{dp}{\frac{\partial f}{\partial x}+p\frac{\partial f}{\partial z}}=\frac{dq}{\frac{\partial f}{\partial y}+q\frac{\partial f}{\partial z}}\\
\frac{\partial f}{\partial x}=xq\\
\frac{\partial f}{\partial y}=q^2\\
\frac{\partial f}{\partial z}=0\\
\frac{\partial f}{\partial p}=xq\\
\frac{\partial f}{\partial q}=xp+2yq"

So, the Chapit's equation is;

"\frac{dp}{pq+p.0}=\frac{dq}{q^2+q.0}=\frac{dz}{-p.xq-q.(xp+2yq)}=\frac{dy}{-(xp+2yp)}=\frac{dx}{-xq}\\
\text{Take the second and fifth fractions}\\
\frac{dq}{q^2}=\frac{dx}{-xq}\\
\frac{dq}{q}+\frac{dx}{x}=0\\

\text{Integrate}\\
\int \frac{dq}{q}+\int \frac{dx}{x}=\int0\\
\ln q+\ln x=\ln a\\
q=\frac{a}{x}\\
\text{Substitute this into (2)}\\
ax^2p+ya^2-x^2=0\\
p=\frac{x^2-ya^2}{ax^2}\\
\text{Substitute the value of p and q into } dz=pdx+qdy\\
dz=\frac{(x^2-a^2y)}{ax^2}dx+\frac{a}{x}dy\\
dz=(\frac{1}{a}-\frac{ay}{x^2})dx+\frac{a}{x}dy\\
\text{Integrate}\\
z=\frac{x}{a}+\frac{ay}{x}+\frac{ay}{x}+b\\
z=\frac{x}{a}+\frac{2ay}{x}+b
 \text{ where a, b are arbitrary constants}"